OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for x in the first equation. x + 3y = 5 x + 7 = 2y What is the resulting equation? 2y+ 2x+ 3y= 5 2y+ 7 + 3y= 5 2y- 7 + 3y= 5

OpenStudy (uri):

I'm too lazy to do it.

OpenStudy (uri):

But no problem,she will help you @xoya.

OpenStudy (anonymous):

The last option because when we substitute it it becomes x + 7 = 2y===> x=2y-7 So when we put the value of x in equation so it becomes x + 3y = 5 ====> 2y-7+3y=5 Got it? :)